Consider two ohmic resistors denoted R1 and R2 which are connected in series with a

cell. Find the resistance and power of R2, given that the current flowing

through R1 and R2 is 0.25 A and that the voltage across the cell is 6 V. R1 = 1Ω.


1
Expert's answer
2021-10-10T16:03:45-0400

The voltage across the combination is


"V=V_1+V_2,\\
V=IR_1+IR_2,\\
IR_2=V-IR_1=6-0.25·1=5.75,\\\space\\
R_2=\frac{5.75}{0.25}
=23\space\Omega."